If z= i log(2-√3) then the value of cos(z) will be
 
  1. i
  2. 2i
  3. 1
  4. 2

Grade:Select Grade

1 Answers

noogler
8 years ago
we know that cosz=eiz+e-iz/2
z=i log (2-root 3)...........iz=-log(2-root3)=log(2+root3)
eiz=2+root3 ….......SUSTITUTE eiz value in cosz
simplfy it, u get cosz=2
 
APPROVE IF USEFULL